OpenStudy (anonymous):

Which of the following is the equation of a circle with center (4, -1) and radius of 8? (x - 4)2 + (y + 1)2 = 64 (x + 4)2 + (y - 1)2 = 64 (x - 4)2 + (y + 1)2 = 8 (x + 4)2 + (y - 1)2 = 8

OpenStudy (stormfire1):

the 3rd one

OpenStudy (anonymous):

The equation of a circle is (x-x1)^2 + (y-y1) = r^2 where (x1,y2) is the center and r is the radius. Plugging 4 in for x, -1 in for y, and 8 in for r, you get: (x-4)^2 + (y+1)^2 = 64, so the third choice is correct.

OpenStudy (anonymous):

the 1st one

OpenStudy (anonymous):

My bad, the first one is correct, sorry.

OpenStudy (stormfire1):

ah yes...the 64!